In the formula V= lwh, find l when w=4in., h=6in., and V =144 cu. in.
Answer by checkley71(8403)

144=L*4*6
144=24L
L=144/24
L=6 IN.
------------------------------------------
TYHE * SIGN MEANS MULTIPLY EXAMPLE:
2*3*4=24